# Where Did This Sentence Come From? Tracing Provenance in LLM Reasoning Distillation

### v1

@@ Abstract
Reasoning distillation, a cost-effective approach for enhancing student model performance, has attracted increasing attention. It typically leverages a large teacher model to generate reasoning paths, which are then used to fine-tune a student model so that it mimics the teacher’s behavior in training contexts. However, previous approaches have lacked a detailed analysis of the origins of the distilled model’s capabilities. It remains unclear whether the student can maintain consistent behaviors with the teacher in novel test-time contexts, or whether it regresses to its original output patterns, raising concerns about the generalization of distillation models. To analyse this question, we introduce a cross-model Reasoning Distillation Provenance Tracing framework. For each action (e.g., a sentence) produced by the distilled model, we obtain the predictive probabilities assigned by the teacher, the original student, and the distilled model under the same context. By comparing these probabilities, we classify each action into four categories: (i) teacher-originated actions, (ii) student-originated actions, (iii) pre-existing actions in both models not enhanced by distillation, and (iv) pre-existing actions boosted through distillation. By systematically disentangling the provenance of each action, we experimentally demonstrate that, in test-time contexts, the distilled model can indeed generate teacher-originated actions, which correlate with and plausibly explain observed performance on distilled model. Building on this analysis, we further propose a teacher-guided data selection method. Unlike prior approach that rely on heuristics (e.g., selecting data most aligned with the student’s original distribution), our method directly compares teacher–student divergences on the training data, providing a principled selection criterion. We validate the effectiveness of our approach across multiple representative teacher models (Deepseek, QwQ, GPT-OSS-120B) and diverse student models (Qwen2.5-7B-Instruct, Qwen4-4B-Base, Qwen3-8B-Base, Qwen3-4B-Instruct-2507). The results highlight the utility of our provenance-tracing framework and underscore its promise for reasoning distillation. We hope to share Reasoning Distillation Provenance Tracing, along with our insights into reasoning distillation, with the community.
@@ 1 Introduction
The rapid development of reinforcement learning (RL) techniques (ppo; grpo; rloo) and resulting large-scale reasoning models (deepseek_r1_0528; gptoss; qwq32b) has accelerated the growth of distillation researches, especially reasoning distillation (am_team_dataset; OpenThoughts). Early work on reasoning distillation focused on creating high-quality open-source datasets (am_team_dataset; OpenThoughts; nvidia_nemotron_nano2), and recent studies have focused on curating and filtering distillation samples to improve efficiency and performance (grape; scar). However, these approaches primarily focus on model performance and fail to provide an explanatory analysis of the sources of the model’s output in test-time contexts. As a result, it remains unclear whether the student model has successfully inherited the knowledge and reasoning logic from the teacher model, raising concerns about the generalization of distillation models (hinton_distilling).
@@
Figure 1: Motivation. In reasoning distillation, the student model learns in the training stage to produce actions consistent with the teacher model within the training context. However, at test time, it remains unclear whether the distilled model will continue to output actions aligned with the teacher model instead of degrading to the outputs of the student model, raising concerns about the generalization of distillation models (hinton_distilling).
To analyse the above problem, we propose a cross-model Reasoning Distillation Provenance Tracing framework. Concretely, we evaluate three open-source distilled models: DeepSeek-Distill-Qwen-7B (deepseek_r1_0528), DeepSeek-R1-0528-Qwen3-8B (deepseek_r1_0528), and LIMO-v2 model (limov2). To analyze the patterns in distilled models’ outputs under test contexts, we collect multiple responses from them on GPQA-D (gpqa) and AIME24. Each response is re-input into the distilled model, the corresponding teacher model, and the original student model. This procedure allows us to obtain, under the same test scenario, the probability assigned by each model to the next action produced by the distilled model. By comparing these probabilities, we can naturally trace the provenance of every action. For example, if the teacher model assigns significantly higher probability to an action than the original student model does, we can attribute that action to teacher-originated actions, since the distilled model’s ability to produce it mainly derives from teacher model’s knowledge transferred during distillation. For further details about other action types, see Section 3.
To this end, we observe that distilled models can reproduce teacher-originated actions in new test contexts. Moreover, these actions are correlated with correct responses on the test set, which helps explain the generalization gains achieved through distillation. This observation further motivates our training design: we hypothesize that when the training data contains a higher proportion of teacher-originated actions, the distilled model attains better final performance. To validate this hypothesis, we propose a teacher-guided data selection strategy that compares the probabilities assigned by the teacher and student models on the training data and prioritizes examples that contain more teacher-originated actions. We then evaluate its effectiveness across multiple settings, including representative teacher models from different families (DeepSeek-R1-671B (deepseek_r1_0528), QwQ-32B (qwq32b), GPT-OSS-120B (gptoss)) and various student models (Qwen2.5-7B-Instruct, Qwen4-4B-Base, Qwen3-8B-Base, Qwen3-4B-Instruct-2507).
@@ 2.1 Reasoning Distillation
Distilling the reasoning abilities of large reasoning models has been an important problem since their emergence (seq_distill; distill_cap2; seq_distill1; seq_distill2). DeepSeek (deepseek_r1_0528) pioneered this line of work by showing that supervised fine-tuning on the outputs of a reasoning teacher, which is also the approach we focus on in this paper, can dramatically enhance the reasoning abilities of smaller models. Numerous subsequent projects (e.g., HuggingFace OpenR1 (openr1), OpenThoughts (OpenThoughts), a-m-team (am_team_dataset), NVIDIA AceReason (nvidia_dataset), Alibaba OmniThought (OpenThoughts), LIMO (limov2), Tencent DeepMath (deepmath)) have devoted substantial effort to constructing large-scale corpora of challenging reasoning problems paired with teacher responses, using rigorous quality filtering, correctness checks, and diversity-aware curation. Most recently, GRAPE (grape) preferentially selects examples whose likelihoods best match the student’s current distribution, thereby steering training toward data that is already well aligned with the student. Rather than focusing solely on artificially designed rules and heuristic rules, we aim to quantify the sources of a distilled model’s capabilities and introduce a data selection criterion that focuses on sentences whose probabilities indicate stronger teacher-originated behavior. This provenance-aware criterion complements prior student-only selection and provides an explicit cross-model signal for reasoning transfer.
@@ 2.2 Model Auditing
Another closely related area is model auditing, a growing line of work that studies (model_audit1; model_audit2; model_audit3) auditing generative models to understand what data they memorize and to attribute outputs back to underlying data sources. In contrast, our work targets model-level provenance in a distillation setting: rather than asking whether specific data are memorized, we aim to trace which upstream models are the sources of a given output, shifting the focus from data membership to the lineage of the models themselves.
@@ 3.3 Analysis on open-source models
Figure 4: Analysis results on LIMO-v2 model (limov2). We apply Reasoning Distillation Provenance Tracing to LIMO-v2 model, analyzing the probability of producing different actions at each action position on two reasoning benchmarks (AIME24 and GPQA-D). Since LIMO-v2 model uses multiple teacher models, we tried different teacher models on all samples. The x-axis denotes the action index, and the y-axis shows the probability of producing that action.
@@ 4.1 Teacher-Guided Data Selection
Figure 5: Comparison between teacher-guided data selection and the previous method (grape).
@@
As shown in Figure 5, the previous method (grape) feeds the training data into the student model before training, computes heuristics metrics based on the logits of student model, and selects samples that more aligned with the student model’s original distribution. In contrast, our proposed method provides a clearer objective for data selection: prioritizing samples where the teacher and student models differ the most.
Moreover, our method introduces only an acceptable additional cost in practice. (1) When one wishes to sample responses, similar to the previous method (grape), our approach incurs no extra overhead, since logits can be obtained directly during sampling. (2) When one wishes to leverage existing open-source distillation datasets for filtering, some additional cost is required. However, it is important to note that re-feeding the generated sequences into the model to extract logits only requires a single forward pass, which is significantly faster than token-by-token generation. Given sufficient GPU memory, we consider the additional time cost to be acceptable.
@@ 4.2 Experiments
We aim to evaluate the effectiveness of teacher-guided data selection across diverse training settings. To this end, we conduct experiments with three distinct teacher models (Deepseek-R1-671B, QwQ-32B, and GPT-OSS-120B) and four student models: Qwen3-4B-Base, Qwen3-8B-Base, Qwen2.5-7B-Instruct, and Qwen3-4B-Instruct-2507. We use two high-quality reasoning datasets, AceReason-1.1-SFT (nvidia_dataset) and OpenThought3-1.2M (OpenThoughts). Each question in these datasets is paired with multiple candidate answers sampled from the teacher model. Further details are provided in Appendix A.5. For each question, we then select a single response using one of three strategies: (i) random selection, which performs similarly to vanilla reasoning distillation (”Vanilla”); (ii) GRAPE (grape) (“GRAPE”); and (iii) our method (“Ours”). All strategies yield training sets of identical size.
For evaluation, we select four widely adopted benchmarks known for their challenging reasoning demands: AIME24, AIME25, MATH500 (math500), and OlympiadBench (olympiadbench). We report the accuracy results.
@@ 4.2.2 Ablation experiments
(3) Effects in different domain. To evaluate the effectiveness of our method in different domain, we used GPT-OSS-120B to sample 10k questions from OpenScienceReasoning_2 (OpenScienceReasoning_2), generating nine candidate responses per question. We then selected a single response for training using different selection strategies. Results are reported in Table 4. We find that our method is effective in the scientific domain and, compared to GRAPE, exhibits stronger generalization in math domain. Analysis on the test set further indicates that the Teacher Sentence in science shares commonalities with that in mathematics, which we hypothesize helps explain why training on scientific data can improve performance on the math test set. Additional analyses are provided in Appendix A.3.3.
@@ A.3 More analysis results
As shown in Figure 3, within DeepSeek-Distill-Qwen-7B, once the sequence length exceeds approximately 4K tokens, the generated outputs are quickly transformed into Teacher Sentence. In contrast, this behavior is not observed in DeepSeek-R1-0528-Qwen3-8B or LIMO-v2 model. We attribute this discrepancy to differences in the underlying student model: DeepSeek-Distill-Qwen-7B is distilled from Qwen2.5-Math-7B (qwen2_5), which has an effective context length of approximately 4K tokens. Beyond this limit, the model’s outputs predominantly reflect patterns inherited from the teacher model. By comparison, DeepSeek-R1-0528-Qwen3-8B and LIMO-v2 model are based on Qwen3-8B-Base (qwen3) and Qwen2.5-32B-Instruct, respectively. These models support substantially longer effective context lengths and thus avoid this limitation.
@@ A.3.2 What behaviors are included in different action types?
We categorize each action using SEAL (seal), which defines three types of behaviors: execution, reflection, and transition. Execution refers to steps that directly advance problem solving, reflection denotes verification, checking, or questioning of the existing reasoning process, and transition represents an intentional change in the current reasoning direction or strategy. Special tokens are excluded from the statistics. The results are shown in Figure 10, from which we make the following three observations.
(1) Both student models exhibit a certain degree of reflection and transition, and reasoning distillation further activates and strengthens these behaviors. Specifically, Student Sentences already contain reflection and transition behaviors, while Boosted Sentences display more of these behaviors than Student Sentences and Shared Sentences after distillation. Although there remains a gap between RL and reasoning distillation, this observation is similar with prior work (qwen25_have_pattern1; qwen25_have_pattern2), which suggests that reflection and transition abilities are already latent in student models and that RL/Reasoning distillation as a post-training method serves to activate them. (2) The vast majority of Shared Sentences are Execution. (3) The student model of Deepseek-R1-0528-Qwen3-8B (Ds-8B) shows stronger reflective ability than the student model of Deepseek-Distill-Qwen-7B (Ds-7B), as evidenced by the higher proportion of reflection behaviors in Student Sentences.
@@ A.4.1 A New Perspective on Understanding Models
Most existing approaches semistic1; semistic2 to understanding model outputs operate in a human-centric semantic space. They typically seek to decompose intermediate activations or outputs into human-interpretable units, such as natural language concepts, symbolic structures, or predefined semantic categories. While this line of work has yielded valuable insights, it implicitly assumes that model behavior is best understood by mapping it onto human-designed semantic structures.
@@ A.6.1 LLM Distillation
Knowledge Distillation. Knowledge distillation was first introduced by (classic_distill1) as a technique for transferring the dark knowledge of a teacher model to a lightweight student model via soft labels, thereby enabling substantial model compression while preserving most of the original accuracy. Subsequent works (classic_distill2; classic_distill3) has extended this framework from multiple perspectives, for example by distilling feature layers or intermediate representations and by introducing relational or structural distillation, so that the student not only mimics the output distribution but also aligns the structural properties of the hidden representation space. Later works (seq_distill1; seq_distill2) further improve alignment by approximating sequence-level KL divergence, leading to more effective distillation. In practice, knowledge distillation (classic_distill4) has been widely applied to tasks such as image classification, object detection, and natural language processing, and has become one of the mainstream approaches for model compression and acceleration.
@@
Reasoning Distillation. Distilling the reasoning abilities of large language models has been an important problem since their emergence (seq_distill; distill_cap2; seq_distill1; seq_distill2). Before the advent of large-scale reasoning models such as O1 (openai_o1), traditional reasoning distillation methods primarily transferred capabilities by aligning intermediate features or output probabilities between teacher and student models (seq_distill; distill_cap2; seq_distill1; seq_distill2). To teach reasoning more explicitly, prior work (distill_cap1; seq_distill; distill_cap2; seq_distill1; seq_distill2) constructs responses that include detailed reasoning traces and trains the student on these signals, thereby strengthening its mastery of reasoning. In the era of large-scale reasoning models such as R1 (deepseek_r1_0528) and QwQ (qwq32b), which naturally exhibit chain-of-thought reasoning and achieve strong performance, distilling their capabilities into smaller models has become an effective and practical path toward improved efficiency. DeepSeek (deepseek_r1_0528) pioneered this line of work by showing that supervised fine-tuning on the outputs of a reasoning teacher, which is also the approach we focus on in this paper, can dramatically enhance the reasoning abilities of smaller models. Numerous subsequent projects (e.g., OpenR1 (openr1), OpenThoughts (OpenThoughts), a-m-team (am_team_dataset), NVIDIA AceReason (nvidia_dataset), OmniThought (OpenThoughts), LIMO (limov2), DeepMath (deepmath)) have devoted substantial effort to constructing and refining large-scale corpora of challenging reasoning problems paired with teacher responses, using rigorous quality filtering, correctness checks, and diversity-aware curation. Most recently, GRAPE (grape) scores candidate responses with the student model and preferentially selects examples whose likelihoods best match the student’s current distribution, thereby steering training toward data that is already well aligned with the student. Rather than focusing solely on artificially designed rules and heuristic rules, we view reasoning distillation as a capability-transfer problem from teacher to student. We aim to quantify the sources of a distilled model’s capabilities: given a context, which actions in a trajectory from distilled model are more likely to originate from the teacher’s behavior rather than the student’s existing tendencies? Building on this perspective, we introduce a data selection criterion that jointly compares teacher–student output distributions and focuses on sentences whose probabilities indicate stronger teacher-originated behavior. This provenance-aware criterion complements prior student-only selection in the following way: it provides an explicit cross-model signal for reasoning transfer. In Section 4, we show that provenance-aware selection outperforms student-only alignment in our settings.

### v2

@@ Abstract
Reasoning distillation, a cost-effective approach for enhancing student model performance, has attracted increasing attention. It typically leverages a large teacher model to generate reasoning paths, which are then used to fine-tune a student model so that it mimics the teacher’s behavior in training contexts. However, previous approaches have lacked a detailed analysis of the origins of the distilled model’s capabilities. It remains unclear whether the student can maintain consistent behaviors with the teacher in novel test-time contexts, or whether it regresses to its original output patterns, raising concerns about the generalization of distillation models. To analyse this question, we introduce a cross-model Reasoning Distillation Provenance Tracing framework. For each action (e.g., a sentence) produced by the distilled model, we obtain the predictive probabilities assigned by the teacher, the original student, and the distilled model under the same context. By comparing these probabilities, we classify each action into four categories: (i) teacher-originated actions, (ii) student-originated actions, (iii) pre-existing actions in both models not enhanced by distillation, and (iv) pre-existing actions boosted through distillation. By systematically disentangling the provenance of each action, we experimentally demonstrate that, in test-time contexts, the distilled model can indeed generate teacher-originated actions, which correlate with and plausibly explain observed performance on distilled model. Building on this analysis, we further propose a teacher-guided data selection method. Unlike prior approach that rely on heuristics (e.g., selecting data most aligned with the student’s original distribution), our method directly compares teacher–student divergences on the training data, providing a principled selection criterion. We validate the effectiveness of our approach across multiple representative teacher models (Deepseek, QwQ, GPT-OSS-120B) and diverse student models (Qwen2.5-7B-Instruct, Qwen3-4B-Base, Qwen3-8B-Base, Qwen3-4B-Instruct-2507). The results highlight the utility of our provenance-tracing framework and underscore its promise for reasoning distillation. We hope to share Reasoning Distillation Provenance Tracing, along with our insights into reasoning distillation, with the community.
footnotetext: Corresponding author
@@ 1 Introduction
The rapid development of reinforcement learning (RL) techniques (Schulman et al., 2017; Shao et al., 2024; Ahmadian et al., 2024) and resulting large-scale reasoning models (DeepSeek-AI, 2025; OpenAI, 2025; Team, 2025b) has accelerated the growth of distillation researches, especially reasoning distillation (Zhao et al., 2025; Guha et al., 2025). Early work on reasoning distillation focused on creating high-quality open-source datasets (Zhao et al., 2025; Guha et al., 2025; NVIDIA, 2025), and recent studies have focused on curating and filtering distillation samples to improve efficiency and performance (Zhang et al., 2025; Li et al., 2025). However, these approaches primarily focus on model performance and fail to provide an explanatory analysis of the sources of the model’s output in test-time contexts. As a result, it remains unclear whether the student model has successfully inherited the knowledge and reasoning logic from the teacher model, raising concerns about the generalization of distillation models (Hinton et al., 2015b).
@@
Figure 1: Motivation. In reasoning distillation, the student model learns in the training stage to produce actions consistent with the teacher model within the training context. However, at test time, it remains unclear whether the distilled model will continue to output actions aligned with the teacher model instead of degrading to the outputs of the student model, raising concerns about the generalization of distillation models (Hinton et al., 2015b).
To analyse the above problem, we propose a cross-model Reasoning Distillation Provenance Tracing framework. Concretely, we evaluate three open-source distilled models: DeepSeek-Distill-Qwen-7B (DeepSeek-AI, 2025), DeepSeek-R1-0528-Qwen3-8B (DeepSeek-AI, 2025), and LIMO-v2 model (Ye et al., 2025). To analyze the patterns in distilled models’ outputs under test contexts, we collect multiple responses from them on GPQA-D (Rein et al., 2024) and AIME24. Each response is re-input into the distilled model, the corresponding teacher model, and the original student model. This procedure allows us to obtain, under the same test scenario, the probability assigned by each model to the next action produced by the distilled model. By comparing these probabilities, we can naturally trace the provenance of every action. For example, if the teacher model assigns significantly higher probability to an action than the original student model does, we can attribute that action to teacher-originated actions, since the distilled model’s ability to produce it mainly derives from teacher model’s knowledge transferred during distillation. For further details about other action types, see Section 3.
To this end, we observe that distilled models can reproduce teacher-originated actions in new test contexts. Moreover, these actions are correlated with correct responses on the test set, which helps explain the generalization gains achieved through distillation. This observation further motivates our training design: we hypothesize that when the training data contains a higher proportion of teacher-originated actions, the distilled model attains better final performance. To validate this hypothesis, we propose a teacher-guided data selection strategy that compares the probabilities assigned by the teacher and student models on the training data and prioritizes examples that contain more teacher-originated actions. We then evaluate its effectiveness across multiple settings, including representative teacher models from different families (DeepSeek-R1-671B (DeepSeek-AI, 2025), QwQ-32B (Team, 2025b), GPT-OSS-120B (OpenAI, 2025)) and various student models (Qwen2.5-7B-Instruct, Qwen3-4B-Base, Qwen3-8B-Base, Qwen3-4B-Instruct-2507).
@@ 2.1 Reasoning Distillation
Distilling the reasoning abilities of large reasoning models has been an important problem since their emergence (Kim and Rush, 2016; Ho et al., 2023; Ko et al., 2024; 2025). DeepSeek (DeepSeek-AI, 2025) pioneered this line of work by showing that supervised fine-tuning on the outputs of a reasoning teacher, which is also the approach we focus on in this paper, can dramatically enhance the reasoning abilities of smaller models. Numerous subsequent projects (e.g., HuggingFace OpenR1 (Hugging Face, 2025), OpenThoughts (Guha et al., 2025), a-m-team (Zhao et al., 2025), NVIDIA AceReason (Liu et al., 2025b), Alibaba OmniThought (Guha et al., 2025), LIMO (Ye et al., 2025), Tencent DeepMath (He et al., 2025)) have devoted substantial effort to constructing large-scale corpora of challenging reasoning problems paired with teacher responses, using rigorous quality filtering, correctness checks, and diversity-aware curation. Most recently, GRAPE (Zhang et al., 2025) preferentially selects examples whose likelihoods best match the student’s current distribution, thereby steering training toward data that is already well aligned with the student. Rather than focusing solely on artificially designed rules and heuristic rules, we aim to quantify the sources of a distilled model’s capabilities and introduce a data selection criterion that focuses on sentences whose probabilities indicate stronger teacher-originated behavior. This provenance-aware criterion complements prior student-only selection and provides an explicit cross-model signal for reasoning transfer.
@@ 2.2 Model Auditing
Another closely related area is model auditing, a growing line of work that studies (Song and Shmatikov, 2019; Carlini et al., 2022; Dekoninck et al., 2024) auditing generative models to understand what data they memorize and to attribute outputs back to underlying data sources. In contrast, our work targets model-level provenance in a distillation setting: rather than asking whether specific data are memorized, we aim to trace which upstream models are the sources of a given output, shifting the focus from data membership to the lineage of the models themselves.
@@ 3.3 Analysis on open-source models
Figure 4: Analysis results on LIMO-v2 model (Ye et al., 2025). We apply Reasoning Distillation Provenance Tracing to LIMO-v2 model, analyzing the probability of producing different actions at each action position on two reasoning benchmarks (AIME24 and GPQA-D). Since LIMO-v2 model uses multiple teacher models, we tried different teacher models on all samples. The x-axis denotes the action index, and the y-axis shows the probability of producing that action.
@@ 4.1 Teacher-Guided Data Selection
Figure 5: Comparison between teacher-guided data selection and the previous method (Zhang et al., 2025).
@@
As shown in Figure 5, the previous method (Zhang et al., 2025) feeds the training data into the student model before training, computes heuristics metrics based on the logits of student model, and selects samples that more aligned with the student model’s original distribution. In contrast, our proposed method provides a clearer objective for data selection: prioritizing samples where the teacher and student models differ the most.
Moreover, our method introduces only an acceptable additional cost in practice. (1) When one wishes to sample responses, similar to the previous method (Zhang et al., 2025), our approach incurs no extra overhead, since logits can be obtained directly during sampling. (2) When one wishes to leverage existing open-source distillation datasets for filtering, some additional cost is required. However, it is important to note that re-feeding the generated sequences into the model to extract logits only requires a single forward pass, which is significantly faster than token-by-token generation. Given sufficient GPU memory, we consider the additional time cost to be acceptable.
@@ 4.2 Experiments
We aim to evaluate the effectiveness of teacher-guided data selection across diverse training settings. To this end, we conduct experiments with three distinct teacher models (Deepseek-R1-671B, QwQ-32B, and GPT-OSS-120B) and four student models: Qwen3-4B-Base, Qwen3-8B-Base, Qwen2.5-7B-Instruct, and Qwen3-4B-Instruct-2507. We use two high-quality reasoning datasets, AceReason-1.1-SFT (Liu et al., 2025b) and OpenThought3-1.2M (Guha et al., 2025). Each question in these datasets is paired with multiple candidate answers sampled from the teacher model. Further details are provided in Appendix A.5. For each question, we then select a single response using one of three strategies: (i) random selection, which performs similarly to vanilla reasoning distillation (”Vanilla”); (ii) GRAPE (Zhang et al., 2025) (“GRAPE”); and (iii) our method (“Ours”). All strategies yield training sets of identical size.
For evaluation, we select four widely adopted benchmarks known for their challenging reasoning demands: AIME24, AIME25, MATH500 (Lightman et al., 2023), and OlympiadBench (He et al., 2024). We report the accuracy results.
@@ 4.2.2 Ablation experiments
(3) Effects in different domain. To evaluate the effectiveness of our method in different domain, we used GPT-OSS-120B to sample 10k questions from OpenScienceReasoning_2 (NVIDIA Corporation, 2024), generating nine candidate responses per question. We then selected a single response for training using different selection strategies. Results are reported in Table 4. We find that our method is effective in the scientific domain and, compared to GRAPE, exhibits stronger generalization in math domain. Analysis on the test set further indicates that the Teacher Sentence in science shares commonalities with that in mathematics, which we hypothesize helps explain why training on scientific data can improve performance on the math test set. Additional analyses are provided in Appendix A.3.3.
@@ A.3 More analysis results
As shown in Figure 3, within DeepSeek-Distill-Qwen-7B, once the sequence length exceeds approximately 4K tokens, the generated outputs are quickly transformed into Teacher Sentence. In contrast, this behavior is not observed in DeepSeek-R1-0528-Qwen3-8B or LIMO-v2 model. We attribute this discrepancy to differences in the underlying student model: DeepSeek-Distill-Qwen-7B is distilled from Qwen2.5-Math-7B (Yang et al., 2024), which has an effective context length of approximately 4K tokens. Beyond this limit, the model’s outputs predominantly reflect patterns inherited from the teacher model. By comparison, DeepSeek-R1-0528-Qwen3-8B and LIMO-v2 model are based on Qwen3-8B-Base (Team, 2025a) and Qwen2.5-32B-Instruct, respectively. These models support substantially longer effective context lengths and thus avoid this limitation.
@@ A.3.2 What behaviors are included in different action types?
We categorize each action using SEAL (Chen et al., 2025), which defines three types of behaviors: execution, reflection, and transition. Execution refers to steps that directly advance problem solving, reflection denotes verification, checking, or questioning of the existing reasoning process, and transition represents an intentional change in the current reasoning direction or strategy. Special tokens are excluded from the statistics. The results are shown in Figure 10, from which we make the following three observations.
(1) Both student models exhibit a certain degree of reflection and transition, and reasoning distillation further activates and strengthens these behaviors. Specifically, Student Sentences already contain reflection and transition behaviors, while Boosted Sentences display more of these behaviors than Student Sentences and Shared Sentences after distillation. Although there remains a gap between RL and reasoning distillation, this observation is similar with prior work (Wang et al., 2025; Gandhi et al., 2025), which suggests that reflection and transition abilities are already latent in student models and that RL/Reasoning distillation as a post-training method serves to activate them. (2) The vast majority of Shared Sentences are Execution. (3) The student model of Deepseek-R1-0528-Qwen3-8B (Ds-8B) shows stronger reflective ability than the student model of Deepseek-Distill-Qwen-7B (Ds-7B), as evidenced by the higher proportion of reflection behaviors in Student Sentences.
@@ A.4.1 A New Perspective on Understanding Models
Most existing approaches Lee et al. (2025); Liu et al. (2025a) to understanding model outputs operate in a human-centric semantic space. They typically seek to decompose intermediate activations or outputs into human-interpretable units, such as natural language concepts, symbolic structures, or predefined semantic categories. While this line of work has yielded valuable insights, it implicitly assumes that model behavior is best understood by mapping it onto human-designed semantic structures.
@@ A.6.1 LLM Distillation
Knowledge Distillation. Knowledge distillation was first introduced by (Hinton et al., 2015a) as a technique for transferring the dark knowledge of a teacher model to a lightweight student model via soft labels, thereby enabling substantial model compression while preserving most of the original accuracy. Subsequent works (SE, 2014; Zagoruyko and Komodakis, 2016) has extended this framework from multiple perspectives, for example by distilling feature layers or intermediate representations and by introducing relational or structural distillation, so that the student not only mimics the output distribution but also aligns the structural properties of the hidden representation space. Later works (Ko et al., 2024; 2025) further improve alignment by approximating sequence-level KL divergence, leading to more effective distillation. In practice, knowledge distillation (Gou et al., 2021) has been widely applied to tasks such as image classification, object detection, and natural language processing, and has become one of the mainstream approaches for model compression and acceleration.
@@
Reasoning Distillation. Distilling the reasoning abilities of large language models has been an important problem since their emergence (Kim and Rush, 2016; Ho et al., 2023; Ko et al., 2024; 2025). Before the advent of large-scale reasoning models such as O1 (Jaech et al., 2024), traditional reasoning distillation methods primarily transferred capabilities by aligning intermediate features or output probabilities between teacher and student models (Kim and Rush, 2016; Ho et al., 2023; Ko et al., 2024; 2025). To teach reasoning more explicitly, prior work (Hsieh et al., 2023; Kim and Rush, 2016; Ho et al., 2023; Ko et al., 2024; 2025) constructs responses that include detailed reasoning traces and trains the student on these signals, thereby strengthening its mastery of reasoning. In the era of large-scale reasoning models such as R1 (DeepSeek-AI, 2025) and QwQ (Team, 2025b), which naturally exhibit chain-of-thought reasoning and achieve strong performance, distilling their capabilities into smaller models has become an effective and practical path toward improved efficiency. DeepSeek (DeepSeek-AI, 2025) pioneered this line of work by showing that supervised fine-tuning on the outputs of a reasoning teacher, which is also the approach we focus on in this paper, can dramatically enhance the reasoning abilities of smaller models. Numerous subsequent projects (e.g., OpenR1 (Hugging Face, 2025), OpenThoughts (Guha et al., 2025), a-m-team (Zhao et al., 2025), NVIDIA AceReason (Liu et al., 2025b), OmniThought (Guha et al., 2025), LIMO (Ye et al., 2025), DeepMath (He et al., 2025)) have devoted substantial effort to constructing and refining large-scale corpora of challenging reasoning problems paired with teacher responses, using rigorous quality filtering, correctness checks, and diversity-aware curation. Most recently, GRAPE (Zhang et al., 2025) scores candidate responses with the student model and preferentially selects examples whose likelihoods best match the student’s current distribution, thereby steering training toward data that is already well aligned with the student. Rather than focusing solely on artificially designed rules and heuristic rules, we view reasoning distillation as a capability-transfer problem from teacher to student. We aim to quantify the sources of a distilled model’s capabilities: given a context, which actions in a trajectory from distilled model are more likely to originate from the teacher’s behavior rather than the student’s existing tendencies? Building on this perspective, we introduce a data selection criterion that jointly compares teacher–student output distributions and focuses on sentences whose probabilities indicate stronger teacher-originated behavior. This provenance-aware criterion complements prior student-only selection in the following way: it provides an explicit cross-model signal for reasoning transfer. In Section 4, we show that provenance-aware selection outperforms student-only alignment in our settings.